OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3.

OpenStudy (anonymous):

y=2*x-3 Y=mx+b

OpenStudy (anonymous):

m is slope and b is y intercept.